# Advanced Geometric Correction Algorithms for 3D Medical Reconstruction: Comparison of Computed Tomography and Macroscopic Imaging

### v1

@@ 1 Introduction
In particular, 3D kidney models obtained via computed tomography offer high accuracy and detail thanks to the use of advanced image reconstruction algorithms. They allow for precise representation of the anatomical and pathological structure of organs. On the other hand, macroscopic models, created based on a series of photographic images of organ sections, can be less accurate due to potential tissue deformations, errors in matching sections, and physiological shrinkage of the extracted organs - for example, due to water loss. The combined use of translation, rotation, and scaling allows effective 3D alignment, compensating for their global effects and improving anatomical consistency. These differences can lead to errors in the interpretation of imaging data, directly affecting the quality of diagnostics and planning of surgical interventions.
@@
By applying advanced image processing techniques and optimization algorithms, it is possible to obtain macroscopic models with significantly better accuracy and anatomical fidelity. Such an approach opens new possibilities in the more accurate modeling of anatomical structures, which is crucial not only in diagnostics and surgery but also in scientific research and medical education.
@@ 2 Problem Statement
The primary issue addressed in this article concerns the discrepancy in geometric accuracy between 3D kidney models obtained from computed tomography (CT) and those derived from macroscopic imaging. These differences can lead to inconsistencies in the interpretation of imaging data, which directly impacts the effectiveness of diagnostics, surgical planning, and medical education in the field of kidney anatomy. This issue is particularly significant in the context of precisely modeling pathological changes, such as kidney tumors, where the accuracy of the model is crucial for treatment success [undefs, undefab, undefx, undefn, undefi, undef].
The challenge lies in developing a method capable of effectively correcting these differences, enabling the creation of accurate and anatomically faithful 3D kidney models based on macroscopic data. Such a method must be able to address potential tissue deformations and errors in matching sections, which are typical for the process of creating macroscopic models. This solution must also be sufficiently flexible to be applied to modeling different types and sizes of kidneys, as well as varying degrees of pathology progression [undefae, undefv].
@@
These differences are not limited to visual aspects but also have a profound impact on the process of creating 3D models. In particular, when constructing 3D kidney models from both CT and macroscopic data, the differences can amount to up to 30% in volume and in specific dimensions, such as length, width, and thickness. Recent comparative studies confirm that macroscopic observation and digital 3D reconstructions (e.g., CT or surface scans) often yield inconsistent measurements and interpretations due to imaging artifacts, handling differences [undefa, undefy]. Such discrepancies pose a significant challenge to the accuracy of anatomical modeling, directly affecting the reliability of the models used in medicine, which underscores the importance of developing an effective method for correcting these differences.
@@
Recent advancements in computational methods and deep learning algorithms have catalyzed a significant evolution in the realm of 3D reconstruction from computed tomography (CT) scans. Contrary to earlier limitations, contemporary research highlights the integration of sophisticated image processing techniques and machine learning models, marking a pivotal enhancement in the precision of anatomical modeling. There are numerous studies on automated CT organ detection systems, which exemplify the integration of advanced computational techniques to improve the accuracy and efficiency of medical diagnostics [undefk, undefg, undefad, undeft, undefb]. This trend is reinforced by comprehensive reviews summarizing the state-of-the-art in 3D medical reconstructions, including both classical and deep learning-based methods [undefag].
@@
The development of CNN-based segmentation for cardiac micro-CT data, for instance, exemplifies the push towards utilizing deep learning for enhancing the granularity and accuracy of 3D reconstructions, underscoring the method’s capability in high-throughput phenotyping applications [undefc]. Similarly, innovations in scaffolding methods for cartilage restoration further illustrate the multifaceted applications of 3D CT reconstructions in regenerative medicine, emphasizing its role in fabricating conducive chondrogenic environments [undefaf].
@@
A critical point in recent research is the correction of geometric artifacts-especially relevant in symmetric organ structures-where sophisticated geometric modeling approaches have been proposed for improving alignment between cross-sectional and full volumetric representations [undefac]. Additionally, learning-based frameworks that incorporate geometric constraints are being explored to optimize model consistency and anatomical fidelity [undefw, undefh].
@@
In contrast to the well-explored domain of CT-based reconstruction, research into macroscopic imaging reconstructions remains underrepresented. Despite the critical insights offered by macroscopic imaging, particularly in surface detail and texture analysis of organs, its potential is yet to be fully leveraged within anatomical modeling [undefu]. The manual delineation of organs, a gold standard in micro-computed tomography (µCT) data processing, highlights the challenges of time-intensive, error-prone methodologies that lack scalability for extensive data analysis [undefu]. Automated segmentation solutions, crucial for processing the extensive field-of-view in CT images, have seen advancements through multi-atlas segmentation (MAS) strategies. However, these often fail to preserve spatial information among adjacent organs, leading to segmented models that may lack coherence [undefr].
@@
Addressing these challenges, recent studies propose innovative approaches for both manual and automated segmentation techniques. For instance, the development of organ-specific insert phantoms through 3D printing technology based on volumetric CT image datasets offers a novel pathway for simulating anatomical features in CT images, providing a cost-effective alternative for organ-specific studies [undef]. Furthermore, the integration of the Internet of Health Things (IoHT) with deep learning for 3D reconstruction presents a forward-thinking solution to obtain high-accuracy diagnosis results, showcasing the potential for super-resolution CT (SRCT) images through advanced computational models [undefae].
@@
Other emerging platforms, such as Nextmed, enable automatic segmentation and visualization of 3D models from imaging data, facilitating reproducible and streamlined workflows in clinical environments [undefj]. At the same time, methods utilizing residual encoder-decoder architectures and directional wavelet transforms have proven effective in reducing noise in low-dose CT reconstructions [undeff, undefp], while deep learning continues to push the frontier of tomographic reconstruction with frameworks designed for image fidelity, generalization, and inference speed [undefz, undefm, undefq, undefaa]. Finally, developments in real-time 3D endoscopic reconstruction methods provide additional pathways for extending anatomical modeling into intraoperative and minimally invasive settings.
@@
The proposed strategy of coupling deterministic optimization with deep learning also aligns with a broader emerging paradigm in computational modeling. Recent studies have demonstrated that hybridizing meta-heuristic or geometric optimization with learning-based ensembles can significantly enhance predictive stability and accuracy in complex structural modeling tasks [undefe]. By leveraging optimization to constrain the learning manifold, these hybrid frameworks overcome the limitations of standalone models, much like our OCM+DL approach for anatomical reconstruction.
Unlike standard DL-only registration, our framework addresses the physical discontinuities and domain gaps inherent in macroscopic imaging. By combining OCM for global topological consistency with DL for local flexibility, we provide a robust solution for cases where fully-learned models typically fail to generalize.
Figure 1: On the left, a CT scan section of a kidney is shown, displaying the organ’s detailed internal structure. On the right, a macroscopic image of a kidney is presented, illustrating its external appearance and shape, providing information on the organ’s surface condition.
Our work integrates slice-wise registration and cross-modal alignment. 3D reconstruction from physical sections is challenged by tissue deformation and ”z-axis” discontinuities, similar to issues in gross pathology. Aligning 2D photography with 3D CT further requires bridging significant domain gaps. Our hybrid framework addresses these by combining deterministic geometric anchors with learned residual refinements for multimodal consistency.
@@ 3 Methods
The proposed methodology employs a two-stage hybrid approach integrating global geometric optimization with local neural refinement. The framework decomposes the registration process to bridge the gap between macroscopic imaging and CT standards, ensuring high data efficiency on small clinical datasets.
@@ 3.1 Data used in the study
The data were collected at the Military Medical Institute in Warsaw, which specializes in the treatment of kidney tumors. This center is equipped with modern CT imaging equipment and experienced staff, ensuring high quality and consistency of the collected data.
Characteristics of kidney tumors in patients: The study included patients diagnosed with various types of kidney tumors or other kidney diseases. Data analysis revealed the following characteristics: Most cases were clear cell renal carcinoma, which is consistent with the general distribution of this type of tumor. The classification of cancers according to the WHO/ISUP (International Society of Urological Pathology) enables determining the tumor’s grade of advancement. Tumors varied in size, indicating their diverse stages of development at the time of diagnosis. For instance, the dimensions of one of the kidneys were $12.5\times 8.5\times 7.5$ cm, while others had smaller dimensions, such as $11\times 4.5\times 7$ cm. Tumors were located in both the right and left kidneys. Provided histopathological descriptions offer information on the tumor’s stage of advancement. For example, one description indicates clear cell renal carcinoma, WHO/ISUP G-4, denoting a high degree of malignancy. These characteristics underscore the diversity of kidney tumor cases in the study group. This diversity significantly impacts the treatment process, as well as the accuracy and efficiency of using CT imaging and macroscopic analysis in diagnostics and surgical planning. The tumor descriptions and classifications are detailed and based on actual clinical data.
Quality and parameters of CT scans: The images used in the study, recorded in DICOM standard version 3, are characterized by a resolution of $512\times 512$ pixels and a 16-bit depth, offering a rich grayscale range essential for precise identification of tissue density differences. Grayscale images are standard for computed tomography, allowing for detailed analysis of anatomical structures. The equipment used in our studies is ’GE MEDICAL SYSTEMS Discovery CT750 HD’. The study focused on detailed analysis of the abdominal cavity and pelvis, conducted in two phases, which is typical for studies requiring high precision imaging.
The layer thickness, measuring 660.6363 micrometers, highlights the high precision of scanning, enabling the detection of minute structures. The kilovolt peak (KVP) set to 120 influences penetration and image contrast, while the exposure time of 6676 milliseconds determines the scanning duration. These technical parameters play a crucial role in obtaining high-resolution and contrast images, essential for accurate diagnostics. The slice thickness of 2.5 mm represents an optimal balance between accuracy and scanning time and radiation dose for the patient. Thinner slices allow for more detailed analysis and reconstruction.
@@
Method of matching scans: Macroscopic images, such as the provided kidney photo after excision, were taken using a Canon EOS 80D DSLR camera under controlled lighting conditions to ensure uniform presentation of tissues and their colors. The camera was equipped with a 50 mm lens at f/2.8 aperture and a 1/60-second exposure time with ISO 200, specifically chosen to capture detailed images crucial for the study’s accuracy and clarity. Each kidney slice was meticulously placed on a standardized background featuring clear scale markers, facilitating later accurate digitization and scaling of images for 3D modeling.
The images were saved in JPEG format, a standard digital format and a resolution of 6000 pixels by 4000 pixels. To minimize the appearance of natural postoperative changes such as minor deformations or color shifts, specific photography protocols were adopted. These protocols, combined with the use of appropriate depth of field and high-quality photographic equipment, ensured the acquisition of sharp and detailed images. The careful setting of these parameters is pivotal to ensuring that the imaging technique used is reproducible in further studies.
@@ 3.2 Implementation Details and Computational Efficiency
The deep learning refinement is performed by a 2D U-Net with a depth of 4 resolution levels. The feature maps follow a $16,32,64,128$ filter sequence, resulting in approximately 1.2 million trainable parameters. This lightweight architecture was chosen to ensure high generalization despite the limited dataset size. Training was conducted for 100 epochs using the Adam optimizer with a learning rate of $10^{-4}$ and a batch size of 8. The loss function combined Binary Cross-Entropy and Dice Loss to ensure both pixel-wise accuracy and regional overlap. All computations were executed on a high-performance server equipped with an NVIDIA GPU (33.70 GB VRAM, Compute Capability 12.0) running CUDA 12.2. A critical advantage of the proposed hybrid framework is its computational profile, as summarized in Table 1. While the deterministic OCM reconstruction serves as the primary bottleneck due to its iterative optimization (taking 12–15 s per slice), the U-Net inference is nearly instantaneous on the GPU (<<0.1 s). This allows the entire 3D kidney reconstruction to be completed in approximately 3 minutes, making it viable for time-sensitive clinical or educational assessments.

### v2

# Deep learning Based Correction Algorithms for 3D Medical Reconstruction in Computed Tomography and Macroscopic Imaging

@@ 1 Introduction
In particular, 3D kidney models obtained via computed tomography offer high accuracy and detail thanks to the use of advanced image reconstruction algorithms. They allow for precise representation of the anatomical and pathological structure of organs. On the other hand, macroscopic models, created based on a series of photographic images of organ sections, can be less accurate due to potential tissue deformations, errors in matching sections, and physiological shrinkage of the extracted organs—for example, due to water loss. The combined use of translation, rotation, and scaling allows effective 3D alignment, compensating for their global effects and improving anatomical consistency. These differences can lead to errors in the interpretation of imaging data, directly affecting the quality of diagnostics and planning of surgical interventions.
@@
By applying advanced image processing techniques and optimization algorithms, it is possible to obtain macroscopic models with significantly better accuracy and anatomical fidelity. Such an approach opens up new possibilities in the more accurate modeling of anatomical structures, which is crucial not only in diagnostics and surgery but also in scientific research and medical education.
@@ 2 Problem Statement
The primary issue addressed in this article concerns the discrepancy in geometric accuracy between 3D kidney models obtained from computed tomography (CT) and those derived from macroscopic imaging. These differences can lead to inconsistencies in the interpretation of imaging data, which directly impact the effectiveness of diagnostics, surgical planning, and medical education in the field of kidney anatomy. This issue is particularly significant in the context of precisely modeling pathological changes, such as kidney tumors, where the accuracy of the model is crucial for treatment success [Pennarossa et al.(2021)Pennarossa, Arcuri, De Iorio, and Gandolfi, Wang et al.(2017)Wang, Ho, Zhang, and Wang, Soler et al.(2014)Soler, Nicolau, Pessaux, and Mutter, Hu et al.(2007)Hu, Penney, Edwards, and Figl, Geiger(1993), Abdullah et al.(2018)Abdullah, McEntee, Reed, and Kench].
The challenge lies in developing a method capable of effectively correcting these differences, enabling the creation of accurate and anatomically faithful 3D kidney models based on macroscopic data. Such a method must be able to address potential tissue deformations and errors in matching sections, which are typical for the process of creating macroscopic models. This solution must also be sufficiently flexible to be applied to modeling different types and sizes of kidneys, as well as varying degrees of pathology progression [Zhang et al.(2020)Zhang, Gong, Yu, Qi, Wen, Hua, et al., Sarmah et al.(2023)Sarmah, Neelima, and Singh].
@@
Figure 1: On the (left), a CT scan section of a kidney is shown, displaying the organ’s detailed internal structure. On the (right), a macroscopic image of a kidney is presented, illustrating its external appearance and shape, providing information on the organ’s surface condition.
These differences are not limited to visual aspects but also have a profound impact on the process of creating 3D models. In particular, when constructing 3D kidney models from both CT and macroscopic data, the differences can amount to up to 30% in volume and in specific dimensions, such as length, width, and thickness. Recent comparative studies confirm that macroscopic observation and digital 3D reconstructions (e.g., CT or surface scans) often yield inconsistent measurements and interpretations due to imaging artifacts, handling differences [Abegg et al.(2021)Abegg, Balbo, Dominguez, Grabherr, Campana, and Moghaddam, Viero et al.(2021)Viero, Biehler-Gomez, Cappella, Messina, Montisci, and Cattaneo]. Such discrepancies pose a significant challenge to the accuracy of anatomical modeling, directly affecting the reliability of the models used in medicine, which underscores the importance of developing an effective method for correcting these differences.
@@
Recent advancements in computational methods and deep learning algorithms have catalyzed a significant evolution in the realm of 3D reconstruction from computed tomography (CT) scans. Contrary to earlier limitations, contemporary research highlights the integration of sophisticated image processing techniques and machine learning models, marking a pivotal enhancement in the precision of anatomical modeling. There are numerous studies on automated CT organ detection systems, which exemplify the integration of advanced computational techniques to improve the accuracy and efficiency of medical diagnostics [Hagen et al.(2021)Hagen, Mair, Bitzer, Bösmüller, and Horger, Chun et al.(2022)Chun, Choi, Kim, Ahn, and Kim, Wang et al.(2018)Wang, Zhang, Jiao, Gao, and Zou, Rister et al.(2020)Rister, Yi, Shivakumar, Nobashi, and Rubin, Armato et al.(2003)Armato, Altman, and La Rivière]. This trend is reinforced by comprehensive reviews summarizing the state of the art in 3D medical reconstructions, including both classical and deep learning-based methods [Zhou et al.(2024)Zhou, Wu, Zuo, Chen, and Hu].
@@
The development of CNN-based segmentation for cardiac micro-CT data, for instance, exemplifies the push towards utilizing deep learning for enhancing the granularity and accuracy of 3D reconstructions, underscoring the method’s capability in high-throughput phenotyping applications [Badea et al.(2024)Badea, Han, Clark, and Badea]. Similarly, innovations in scaffolding methods for cartilage restoration further illustrate the multifaceted applications of 3D CT reconstructions in regenerative medicine, emphasizing its role in fabricating conducive chondrogenic environments [Zhao et al.(2024)Zhao, Yan, Ding, Dai, Feng, Li, Ma, et al.].
@@
A critical point in recent research is the correction of geometric artifacts—especially relevant in symmetric organ structures—where sophisticated geometric modeling approaches have been proposed for improving alignment between cross-sectional and full volumetric representations [Wang et al.(2024)Wang, Sun, Li, Lin, Wang, and Cui]. Additionally, learning-based frameworks that incorporate geometric constraints are being explored to optimize model consistency and anatomical fidelity [Shen et al.(2022)Shen, Zhao, Capaldi, Pauly, and Xing, Feng et al.(2023)Feng, Yang, Guo, and Li].
@@
In contrast to the well-explored domain of CT-based reconstruction, research into macroscopic imaging reconstructions remains under-represented. Despite the critical insights offered by macroscopic imaging, particularly in surface detail and texture analysis of organs, its potential is yet to be fully leveraged within anatomical modeling [Rosenhain et al.(2018)Rosenhain, Magnuska, Yamoah, Rawashdeh, Kiessling, Gremse, et al.]. The manual delineation of organs, a gold standard in micro-computed tomography (µCT) data processing, highlights the challenges of time-intensive, error-prone methodologies that lack scalability for extensive data analysis [Rosenhain et al.(2018)Rosenhain, Magnuska, Yamoah, Rawashdeh, Kiessling, Gremse, et al.]. Automated segmentation solutions, crucial for processing the extensive field-of-view in CT images, have seen advancements through multi-atlas segmentation (MAS) strategies. However, these often fail to preserve spatial information among adjacent organs, leading to segmented models that may lack coherence [Oliveira et al.(2018)Oliveira, Queirós, Morais, Torres, Gomes-Fonseca, Fonseca, and Vilaça].
@@
Addressing these challenges, recent studies propose innovative approaches for both manual and automated segmentation techniques. For instance, the development of organ-specific insert phantoms through 3D printing technology based on volumetric CT image datasets offers a novel pathway for simulating anatomical features in CT images, providing a cost-effective alternative for organ-specific studies [Abdullah et al.(2018)Abdullah, McEntee, Reed, and Kench]. Furthermore, the integration of the Internet of Health Things (IoHT) with deep learning for 3D reconstruction presents a forward-thinking solution to obtain high-accuracy diagnosis results, showcasing the potential for super-resolution CT (SRCT) images through advanced computational models [Zhang et al.(2020)Zhang, Gong, Yu, Qi, Wen, Hua, et al.].
@@
Other emerging platforms, such as Nextmed, enable automatic segmentation and visualization of 3D models from imaging data, facilitating reproducible and streamlined workflows in clinical environments [González Izard et al.(2020)González Izard, Sánchez Torres, Alonso Plaza, Juanes Mendez, and García-Peñalvo]. At the same time, methods utilizing residual encoder–decoder architectures and directional wavelet transforms have proven effective in reducing noise in low-dose CT reconstructions [Chen et al.(2017)Chen, Zhang, Kalra, Lin, Chen, Liao, Zhou, and Wang, Kang et al.(2017)Kang, Min, and Ye], while deep learning continues to push the frontier of tomographic reconstruction with frameworks designed for image fidelity, generalization, and inference speed [Wang(2016), Hsieh et al.(2013)Hsieh, Nett, Yu, Sauer, Thibault, and Bouman, Kofler et al.(2020)Kofler, Haltmeier, Schaeffter, Kachelrieß, Dewey, Wald, and Kolbitsch, Wang et al.(2018)Wang, Ye, Mueller, and Fessler]. Finally, developments in real-time 3D endoscopic reconstruction methods provide additional pathways for extending anatomical modeling into intraoperative and minimally invasive settings.
@@
The proposed strategy of coupling deterministic optimization with deep learning also aligns with a broader emerging paradigm in computational modeling. Recent studies have demonstrated that hybridizing meta-heuristic or geometric optimization with learning-based ensembles can significantly enhance predictive stability and accuracy in complex structural modeling tasks [Bouhacina et al.(2026)Bouhacina, Khatir, Matoug, and Tamine]. By leveraging optimization to constrain the learning manifold, these hybrid frameworks overcome the limitations of standalone models, much like our OCM + DL approach for anatomical reconstruction.
Unlike standard DL-only registration, our framework addresses the physical discontinuities and domain gaps inherent in macroscopic imaging. By combining OCM for global topological consistency with DL for local flexibility, we provide a robust solution for cases where fully learned models typically fail to generalize.
Our work integrates slice-wise registration and cross-modal alignment. 3D reconstruction from physical sections is challenged by tissue deformation and “z-axis” discontinuities, similar to issues in gross pathology. Aligning 2D photography with 3D CT further requires bridging significant domain gaps. Our hybrid framework addresses these by combining deterministic geometric anchors with learned residual refinements for multimodal consistency.
@@ 3 Methods
blueThe proposed methodology employs a two-stage hybrid approach integrating global geometric optimization with local neural refinement. The framework decomposes the registration process to bridge the gap between macroscopic imaging and CT standards, ensuring high data efficiency on small clinical datasets. The overall architecture of the proposed framework is illustrated in Figure 2.
@@ 3.1 Data Used in the Study
The data were collected at the Military Institute of Medicine-National Research Institute in Warsaw, which specializes in the treatment of kidney tumors. This center is equipped with modern CT imaging equipment and experienced staff, ensuring high quality and consistency of the collected data.
Characteristics of kidney tumors in patients: The study included patients diagnosed with various types of kidney tumors or other kidney diseases. Data analysis revealed the following characteristics: Most cases were clear cell renal carcinoma, which is consistent with the general distribution of this type of tumor. The classification of cancers according to the WHO/ISUP (International Society of Urological Pathology) enables determining the tumor’s grade of advancement. Tumors varied in size, indicating their diverse stages of development at the time of diagnosis. For instance, the dimensions of one of the kidneys were $12.5\times 8.5\times 7.5$ cm, while others had smaller dimensions, such as $11\times 4.5\times 7$ cm. Tumors were located in both the right and left kidneys. The provided histopathological descriptions offer information on the tumor’s stage of advancement. For example, one description indicates clear cell renal carcinoma, WHO/ISUP G-4, with infiltration of the renal vein wall and perinephric fat, denoting a high degree of malignancy. These characteristics underscore the diversity of kidney tumor cases in the study group. This diversity significantly impacts the treatment process, as well as the accuracy and efficiency of using CT imaging and macroscopic analysis in diagnostics and surgical planning. The tumor descriptions and classifications are detailed and based on actual clinical data.
Quality and parameters of CT scans: The images used in the study, recorded in DICOM standard version 3, are characterized by a resolution of $512\times 512$ pixels and a 16-bit depth, offering a rich grayscale range essential for precise identification of tissue density differences. Grayscale images are standard for computed tomography, allowing for detailed analysis of anatomical structures. The equipment used in our studies is blueGE Medical Systems Discovery CT750 HD (GE Healthcare, Chicago, IL, USA). The study focused on detailed analysis of the abdominal cavity and pelvis, conducted in two phases, which is typical for studies requiring high-precision imaging.
The layer thickness, measuring 660 660.6363 micrometers highlights, the high precision of scanning, enabling the detection of minute structures. The kilovolt peak (KVP) set to 120 influences penetration and image contrast, while the exposure time of 6676 milliseconds determines the scanning duration. These technical parameters play a crucial role in obtaining high-resolution and contrast images, essential for accurate diagnostics. The slice thickness of 2.5 mm represents an optimal balance between accuracy and scanning time and radiation dose for the patient. Thinner slices allow for more detailed analysis and reconstruction.
@@
Method of matching scans: Macroscopic images, such as the provided kidney photo after excision, were taken using a blueCanon EOS 80D DSLR camera (Canon Inc., Tokyo, Japan) under controlled lighting conditions to ensure uniform presentation of tissues and their colors. The camera was equipped with a 50 mm lens at f/2.8 aperture and a 1/60 s exposure time with ISO 200, specifically chosen to capture detailed images crucial for the study’s accuracy and clarity. Each kidney slice was meticulously placed on a standardized background featuring clear scale markers, facilitating later accurate digitization and scaling of images for 3D modeling.
The images were saved in JPEG format, a standard digital format, with a resolution of 6000 pixels by 4000 pixels. To minimize the appearance of natural postoperative changes such as minor deformations or color shifts, specific photography protocols were adopted. These protocols, combined with the use of an appropriate depth of field and high-quality photographic equipment, ensured the acquisition of sharp and detailed images. The careful setting of these parameters is pivotal to ensuring that the imaging technique used is reproducible in further studies.
@@ 3.2 Implementation Details and Computational Efficiency
The deep learning refinement was performed by a 2D U-Net with a depth of 4 resolution levels. The feature maps followed a $16,32,64,128$ filter sequence, resulting in approximately 1.2 million trainable parameters. This lightweight architecture was chosen to ensure high generalization despite the limited dataset size. Training was conducted for 100 epochs using the Adam optimizer with a learning rate of $10^{-4}$ and a batch size of 8. The loss function combined Binary Cross-Entropy and Dice Loss to ensure both pixel-wise accuracy and regional overlap. All computations were executed on a high-performance server equipped with an NVIDIA GPU (33.70 GB VRAM, Compute Capability 12.0) running CUDA 12.2. A critical advantage of the proposed hybrid framework is its computational profile, as summarized in Table 1. While the deterministic OCM reconstruction serves as the primary bottleneck due to its iterative optimization (taking 12–15 s per slice), the U-Net inference is nearly instantaneous on the GPU (<<0.1 s). This allows the entire 3D kidney reconstruction to be completed in approximately 3 min, making it viable for time-sensitive clinical or educational assessments.